Suppose milk and honey are the only products produced in Ghana. Use the data belowfrom the

ii.real GDP, (3marks)

a.Computefor each year the

economy of Ghana to answer the questions that follow:

iii.the GDP deflator.(3marks)

YearPrice of MilkQuantity of MilkPrice of HoneyQuantity of Honey

i.nominal GDP, (3marks)

2012Ghc1200Ghc2100

2013Ghc2200Ghc4100

2011Ghc1100Ghc250

Base year: 2011

b.Compute for 2012 and 2013 from the respective preceding years the percentage change in

Required:

i.nominal GDP, (2marks)

ii.real GDP, (2marks)

iii.the GDP deflator (2marks)


1
Expert's answer
2021-08-04T02:43:01-0400

Real GDP

2012:

"(1\times200)+(2\times200)=400"

2013:

"(2\times 200)+(4\times100)=800"

2011:

"(1\times100)(2\times50)=200"

Nominal GDP

2011:

"(1\times100)+(2\times50)=200"

2012:

"(1\times200)+(2\times100)=400"

2013:

"(1\times200)+(2\times100)=400"

Deflator

"=\frac {Nominal GDP}{Real GDP}\times100"


2011: "=\frac{200}{200}\times100=100"%

2012: "=\frac {400}{400}\times100=100"%

2013: "=\frac {400}{800}\times100=50" %

% change in nominal GDP

2011-2012:

"=\frac{400-200}{200}\times100=100"%

2012-2013:

"=\frac {400-400}{400}\times100=0" %

% change in real GDP

2011-2012:

"=\frac{400-200}{200}\times100=100" %

2012-2013:

"=\frac {800-400}{400}\times100=100" %

% change in deflator

2011-2012:"=100-100=0" %

2012-2013:"=50-100=-50" %